# The Overcooked Generalisation Challenge

### v2

@@ 6 Analysing & Benchmarking
Figure 10: Sample levels that our models perform best (top) and worst in (bottom). The number of visits to each grid cell is shown as a heatmap overlay, while the mean return is below. The layouts where the model performs worst tend to feature narrow elements or large distances between items.
We perform two final experiments to investigate our baselines’ poor performance and eliminate trivial sources of errors. First, we hypothesise that the top-down observations in OGC make generalisation difficult as they are not invariant to mirroring or rotations (Ye et al., 2020). To test this, we evaluated agents on 24 hand-designed circular evaluation levels with different kinds of symmetry, as shown in Figure 14. We found that agents tend to perform similarly across these layouts as the standard deviation is at most $1.1$, and therefore reject this hypothesis (more details in Appendix A.7). Second, we investigate the kinds of levels our best-performing model does well vs poorly in from a pool of randomly generated evaluation levels in Figure 10. The figure summarises the cooperation behaviour of the agents by showing which cells are visited most frequently to give an impression of their motion patterns. While on many layouts, our model reaches good self-play performance (up to a maximum mean reward of $84.4$; top row), it typically delivers few to no soups in layouts that it performs worst in. These levels tend to be narrow/convoluted and/or feature big distances between objects. In conclusion, current DCD methods struggle with generating training layouts of correct complexity, i.e. ones that are similarly hard to evaluation ones.
Discussion
Previous work (Jiang et al., 2021a) found that PLR outperforms the other here-tested algorithms in navigation-based tasks. Our more challenging environment suggests that this might not always be the case. In our preliminary analysis, PAIRED outperformed other DCD methods. Compared to mazes, car racing, or walker environments with fewer moving pieces, Overcooked layouts are more complex to design, requiring the designer to place multiple objects in relation to each other and the agents (see Figure 2). Therefore, methods that employ a random generator struggle in such a big design space. This thus requires a capable generator and suggests that simple navigation-based environments used to benchmark DCD in UED algorithms do not allow full performance evaluation. As such, OvercookedUED can be an important part of evaluating DCD algorithms. We envision that general Overcooked agents should be evaluated in scenarios that are difficult for self-play agents using our benchmark. These include coordinating with strongly-biased agents (Yu et al., 2023) in Coordination Ring (see Section 1) and Asymmetric Advantages as described in (Ruhdorfer, 2023) and for which we provide the tools.
Table 2: Performance on mirrored and rotated levels, illustrated in Figure 14. The low variance suggests that the agents do not struggle with spatial reasoning.
Method SoftMoE-LSTM CNN-S5 CNN-LSTM
DR $11.91\pm 0.8$ $0.00\pm 0$ 0.0 $1.96\pm 0.3$
PLR $4.39\pm 0.4$ $0.49\pm 0.2$ $0.65\pm 0.2$
PAIRED $36.83\pm 1.1$ $0.58\pm 0.2$ $0.59\pm 0.1$
ACCEL $2.36\pm 0.3$ $1.69\pm 0.2$ $0.78\pm 0.2$
7 Limitations
Despite its many advantages, our challenge has two main limitations. First, we artificially restricted the maximum size of the layouts to allow the environment. Future work should focus on more natural representations of the whole scene, e.g. using graphs or item embeddings. While we included a partial observation that could theoretically be computed independently of size, similar to the vector-based observation used for behaviour cloning agents in (Carroll et al., 2019), batching across layouts in OvercookedUED still requires the layouts to be scaled to the same height and width. Second, while our challenge allows us to study zero-shot coordination via generalising across layouts, reasoning about other agents (Rabinowitz et al., 2018; Gandhi et al., 2021; Bara et al., 2023; Bortoletto et al., 2024b, a) might be equally important to achieve zero-shot cooperation capabilities on unknown layouts. This is plausible given that humans can reason about the mental states of other agents via Theory of Mind (Premack & Woodruff, 1978), as well as the physical configuration of the space in which they operate. Future work could thus explore reasoning about other agents in previously unexplored environments.
@@ 8 Conclusion
We presented the Overcooked Generalisation Challenge (OGC) – a generalisation challenge focusing on (zero-shot) cooperation in MARL in out-of-distribution test levels. Our challenge is the first open-source cooperative multi-agent UED environment, which is significantly more challenging than previous environments commonly used in UED and DCD research. In addition to using the challenge in UED research, we have shown how the OGC can be used in future research on human-AI collaboration as a zero-shot cooperation benchmark for general agents. That is, our challenge establishes a link between generalisation and zero-shot coordination. Our work is the first to provide the research community with the tools to train and evaluate agents capable of coordinating in previously unknown physical spaces and with novel partners.
Impact Statement
While our work is largely foundational and concerned with providing the research community with the appropriate tools for the training and evaluation of agents in game-like environments, special caution is always imperative should this research be applied to human-AI collaboration. Even though our goal is to improve collaboration, safeguards should be applied to make sure that humans are always safe from harm. Especially in real-world applications where accidents could potentially result in bodily harm. Since our work is still far removed from any real-world application, we do not expect that our work in its present form carries the risk of materialising these harms. Some form of unsupervised environment design in collaborative environments might be part of future systems, and we, therefore, acknowledge these risks. This work, of course, also carries the potential to improve human-AI collaboration, and we make an important contribution to advancing the field with potential impacts in all kinds of human-machine interaction.
@@ A.2 Infrastructure & tools
We ran our experiments on a server running Ubuntu 22.04, equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs. All training runs are executed on a single GPU only. We trained our models using Jax (Bradbury et al., 2018) and Flax (Heek et al., 2023) with 1, 2 and 3 as random seed for training DCD methods and 1 to 8 as random seeds for the populations. Training the DCD methods usually finishes in under 24 hours, only SoftMoE and PAIRED-based methods take longer. SoftMoE-based policies often take an extra 50% wall-clock time to train. Noticeable is also that our S5 implementation is the fastest, usually needing roughly 30% less time. Both are compared to the default architectures’ training time. In the longest case, the combination of a SoftMoE-LSTM policy trained with PAIRED takes about 80 hours to complete training. Our benchmark should be runnable on any system that features a single CUDA-compatible GPU. Although in our experience our experiments will require 32GB VRAM to run.
@@ A.3 Extended Related Work
We present an overview over how our environment compares to other UED environments in Table 3. Additionally, we compare our work to single-agent UED environments and other attempts at using procedural contentment generation for Overcooked-AI below.
Single-Agent UED Environments
Early work on generalisation mainly focused on single-agent environments (Zhang et al., 2018b; Farebrother et al., 2018; Cobbe et al., 2019) and these are also popular in UED research. Among these, prior work has studied mazes (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022; Jiang et al., 2023; Li et al., 2023a; Beukman et al., 2024), bipedal walkers (Wang et al., 2019, 2020; Parker-Holder et al., 2022) or car racing environments (Jiang et al., 2021a). One likely reason for their popularity as benchmarks for DCD is that new levels are easy to generate, and agents are usually fast to train. However, they are limited to a single agent, with limited options to interact with the environment and other agents, and thus bear little resemblance to real-world problems.
Procedual-Content Generation in Overcooked
Related to our work is the work of (Fontaine et al., 2021) in which the authors used procedurally generated Overcooked layouts to evaluate the impact of different layouts on human-robot interaction using planning algorithms. However, while they use procedural content generation in the Overcooked context their research does not focus on cross-layout generalisation – a major theme in our work.
@@ A.4 Hyperparameters
Learning Rate $\eta$ [$1\cdot 10^{-4}$, $\mathbf{3\cdot 10^{-4}}$, $5\cdot 10^{-4}$, $1\cdot 10^{-3}$]
Entropy Coefficient [$\mathbf{0.01}$ $0.1$]
@@
# Hidden Layers [$2$, $\mathbf{3}$, $4$]
Reward Shaping Annealing Steps [00, $2500000$, $5000000$, until end]
@@
$n$ walls to place Sampled between 0015151515
$n$ onion piles to place Sampled uniformly between 11112222
$n$ plate piles to place Sampled uniformly between 11112222
$n$ pots to place Sampled uniformly between 11112222
$n$ goals to place Sampled uniformly between 11112222
@@
$n$ walls to place Sampled between 0015151515
@@
In the experiment displayed in Figure 5 we show how policies behave when trained on all versus on only one layout and then decide on the SoftMoE architecture for our agents. Since these are considerably easier problems we train these models on fewer environment steps in total. We train the overfitting baseline for roughly $1/30$ ($12,800,00$ steps in the environment) of the experience of all UED methods and only use $1,000$ outer loops. For the Oracle baseline, we use $1/2$ of the experience ($200,000,000$ steps in the environment) and only $5,000$ outer loops. Both are trained until convergence and to speed up training we deploy 100 environment simulators. Recall that the UED methods use $30,000$ training loops and $400,000,000$ steps in the environment. Notably, in the case of population PAIRED these steps apply per student and do not include any additional steps taken in the teacher environment.
@@ A.6.2 Performance across levels
Table 12: Mean episode solved rate for the different methods averaged over the respective testing layouts. The best result is shown in bold. We report aggregate statistics over three random seeds. As a baseline we include an Oracle version for all architectures, which was trained on the five testing layouts directly.
@@
DR $0.02\pm 0.0\%$ $6.31\pm 10.1\%$ $0.00\pm 0.0\%$
PLR⟂,∥ $0.00\pm 0.0\%$ $0.33\pm 0.3\%$ $0.00\pm 0.0\%$
Pop. PAIRED $0.13\pm 0.2\%$ $\mathbf{11.46\pm 2.1\%}$ $0.00\pm 0.0\%$
ACCEL $0.00\pm 0.0\%$ $0.00\pm 0.0\%$ $0.00\pm 0.0\%$
Oracle $95.40\pm 7.5\%$ $99.67\pm 0.6\%$ $97.53\pm 4.1\%$
To accompany the overall performance measured by reward in the main paper in Table 1 we also measure the mean solved rate on display it in Table 12.
@@ A.6.3 Performance on individual levels
Table 13: Performance on all evaluation layouts. We show the mean episode reward R and the mean episode solved rate SR. The overall best result per layout is presented in bold excluding oracle results.
@@
Cramped DR 1.70 0.0% 1.54 0.2% 0.00 0.0%
PLR⟂,∥ 1.12 0.0% 5.02 2.1% 0.14 0.0%
Pop. PAIRED 0.8 0.0% 15.33 17.0% 0.00 0.0%
ACCEL 1.93 0.00% 2.53 0.0% 0.46 0.0%
Oracle 241.27 96.7% 245.54 100.0% 189.47 99.7%
Coord DR 0.00 0.0% 0.00 0.0% 0.00 0.0%
@@
Oracle 197.8 100.0% 204.53 100.0% 119.33 99.0%
Forced DR 0.00 0.0% 0.02 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.02 0.0% 0.02 0.0%
Pop. PAIRED 0.00 0.0% 0.00 0.0% 0.00 0.0%
ACCEL 0.00 0.0% 0.67 0.0 % 0 0.0 %
Oracle 196.8 100.0% 204.53 100.0% 133.47 94.7%
Asymm DR 0.58 0.1% 8.64 4.4% 0.00 0.0%
PLR⟂,∥ 0.08 0.0% 0.10 0.0% 0.08 0.0%
Pop. PAIRED 2.4 0.6% 28.67 40.4% 0.00 0.0%
ACCEL 0.67 0.0% 1.00 0.0 % 0.00 0.0 %
Oracle 220.4 100.0% 277.8 98.4% 247.87 99.7%
Counter DR 0.00 0.0% 0.00 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.00 0.0% 0.00 0.0%
Pop. PAIRED 0.00 0.0% 0.14 0.0% 0.00 0.0%
ACCEL 0.00 0.0% 0.00 0.0 % 0.00 0.0 %
Oracle 91.2 77.3% 152.73 100.0% 84.93 94.7%
@@ A.6.4 Population training details
Both populations were trained over 8 random seeds. As architecture, we used a simple CNN encoder without RNN as in prior work (Zhao et al., 2023; Yu et al., 2023). To give an intuition into the performance of the members of the population, we present the training curves over all 8 seeds of training an FCP population in Figure 12. MEP was trained with the same architecture, with the same amount of experience per agent and achieved practically identical results. As in prior work (Zhao et al., 2023) we set the population entropy coefficient during training to $\alpha=0.01$.
@@ A.6.5 Detailed results with populations
DR $18.18\pm 1.69$ $6.86\pm 5.27$ $65.05\pm 5.15$ $1.09\pm 0.21$ $17.88\pm 10.27$
PLR⟂,∥ $7.64\pm 0.89$ $5.60\pm 1.29$ $60.35\pm 6.89$ $1.76\pm 0.86$ $21.90\pm 1.26$
Pop. PAIRED $\mathbf{42.28\pm 19.59}$ $\mathbf{10.12\pm 1.67}$ $\mathbf{63.41\pm 9.13}$ $\mathbf{2.57\pm 1.46}$ $\mathbf{21.97\pm 2.73}$
ACCEL $8.19\pm 1.08$ $9.39\pm 3.21$ $61.67\pm 2.79$ $2.04\pm 2.37$ $17.94\pm 2.29$
@@
DR $19.32\pm 0.39$ $18.04\pm 5.75$ $\mathbf{62.77\pm 7.22}$ $1.69\pm 0.67$ $30.35\pm 4.42$
PLR⟂,∥ $7.53\pm 0.92$ $\mathbf{21.23\pm 1.91}$ $57.2\pm 4.4$ $2.45\pm 1.23$ $2.45\pm 1.23$
Pop. PAIRED $\mathbf{42.66\pm 20.31}$ $18.34\pm 4.85$ $61.64\pm 7.63$ $\mathbf{3.58\pm 1.69}$ $\mathbf{31.24\pm 5.52}$
ACCEL $9.08\pm 1.11$ $18.43\pm 1.77$ $53.02\pm 5.53$ $2.88\pm 3.31$ $28.81\pm 2.01$
@@
DR $24.19\pm 4.60\%$ $4.56\pm 5.32\%$ $72.11\pm 6.29\%$ $0.01\pm 0.01\%$ $23.76\pm 18.85\%$
PLR⟂,∥ $8.84\pm 1.31\%$ $2.04\pm 0.95\%$ $68.14\pm 1.21\%$ $0.11\pm 0.12\%$ $\mathbf{30.89\pm 3.83\%}$
Pop. PAIRED $\mathbf{56.91\pm 25.08\%}$ $6.07\pm 2.54\%$ $\mathbf{72.48\pm 6.14\%}$ $0.2\pm 0.41\%$ $30.16\pm 7.7\%$
ACCEL $8.79\pm 1.59\%$ $\mathbf{6.45\pm 5.18\%}$ $68.09\pm 2.16\%$ $\mathbf{0.51\pm 0.89\%}$ $20.67\pm 5.18\%$
@@
DR $26.34\pm 3.55\%$ $27.41\pm 10.31\%$ $70.78\pm 4.23\%$ $0.05\pm 0.07\%$ $50.07\pm 6.67\%$
PLR⟂,∥ $8.24\pm 1.28\%$ $\mathbf{33.76\pm 4.89\%}$ $65.38\pm 4.55\%$ $0.28\pm 0.41\%$ $\mathbf{50.97\pm 4.01\%}$
Pop. PAIRED $\mathbf{57.43\pm 26.49\%}$ $24.72\pm 10.42\%$ $\mathbf{72.97\pm 7.6\%}$ $0.4\pm 0.5\%$ $50.64\pm 7.5\%$
ACCEL $9.16\pm 2.14\%$ $25.91\pm 4.56\%$ $64.31\pm 4.01\%$ $\mathbf{1.38\pm 2.33\%}$ $49.23\pm 2.89\%$
@@
A.7 Error Analysis
Figure 14: An illustration of the circular evaluation levels; we move the kitchen around the sides and vary the size.
We hypothesise that agents may fail to generalise since observations are hard to generalise from. To test this we design testing layouts that rotate and mirror features of the environment to look for systemic failures, i.e. cases in which an agent does well in one environment but not its mirrored version. While methods generally differed in how well they performed along the same line as previous analysis suggests, the low standard deviations show that any given method performs similarly well on each of the 24 layouts, see Table 2.
A.7.1 Training curves and evaluation

### v3

@@ 5.4 Experiment 4: Qualitative Failure Analysis
Figure 11: Sample levels in which our models perform best and worst. The number of visits to each grid cell is shown as a heatmap overlay, while the mean return is below. The worst layouts tend to feature narrow corridors or large distances between items.
We perform a final experiment to investigate the agents’ poor performance. To better understand which structures impede performance, we show cell visit patterns for the best and worst-performing layouts in Figure 11. While on many layouts, our PAIRED SoftMoE-LSTM agent reached good self-play performance (up to a maximum mean reward of $84.4$; top row), it delivers few to no soups in others. Poorly performing layouts often feature narrow corridors or large object distances, suggesting that agents fail in environments requiring fine-grained coordination and collision avoidance. We previously identified a key failure mode: (1) agents fail to learn spatially invariant coordination strategies. This experiment reveals a second: (2) performance degrades significantly in more complex spatial layouts, particularly those that demand structured movement and proximity-based interaction. These insights underline the difficulty of OGC and the need for curriculum strategies that better expose agents to high-complexity, high-coordination scenarios during training.
6 Discussion
Our results suggest that generalising to both novel partners and novel environments remains a fundamental challenge in cooperative reinforcement learning. While e.g. Jha et al. (2025) explored cross-environment play, a preliminary analysis showed that their agents also do not generalise to entirely novel layouts/partners. To show this, we retrained their agents and tested them on the 32 evaluation layouts discussed in Experiment 2. When evaluated on the 32 transformed layouts from Experiment 2, their method only achieved an average return of $0.46\pm 2.6$, performing worse than the proposed UED-based approaches (e.g., PAIRED with SoftMoE-LSTM). These agents appear to learn brittle strategies tied to the five original Overcooked layouts, and are unable to adapt to structural variations or unfamiliar partners.
Our findings also have direct implications for future research on UED and DCD: While previous studies, such as Jiang et al. (2021a), found that PLR outperformed other DCD methods in navigation-based tasks, we show that this result does not generalise to more complex, multi-agent cooperative environments. In our experiments, PAIRED consistently outperformed PLR and ACCEL. We attribute this to the increased design space complexity offered by the OGC: The environment contains multiple object types, coordination bottlenecks, and sparse rewards, all of which demand deliberate environment construction. In simpler domains like mazes or locomotion tasks, randomly generated or lightly curated curricula may suffice. In OvercookedAI, however, this approach fails to expose agents to the kinds of structured coordination tasks they must solve at test time.
This leads to three key conclusions: First, DCD methods must scale with environment complexity. Benchmarks that rely on narrow design spaces (e.g., only walls in a maze) are insufficient for evaluating the capabilities of curriculum-learning algorithms. The OGC reveals that without principled curriculum generation, agents may never encounter useful learning signals. Second, Current DCD methods do not scale natively to realistic cooperative tasks. Even with tuned architectures and training regimes, we observed poor generalisation across environments and partners. This highlights the importance of new methods that integrate environmental and partner generalisation. Third, the OGC provides a critical testbed for advancing this next generation of methods. By supporting both axes of generalisation, the OGC offers a foundation for future research into UED-ZSC methods capable of producing robust, general-purpose cooperative agents that perform well in open-ended multi-agent settings.
7 Limitations & Future Work
Despite its advantages, we also identified two limitations of the OGC: First, to support parallel training and JAX-based acceleration, we constrain all layouts to a fixed maximum height and width. While we included a partial observation that can theoretically be computed independently of size, similar to the vector-based observation used for behaviour cloning agents in (Carroll et al., 2019), batching across layouts in OvercookedUED still requires the layouts to be scaled to the same height and width. Future work could explore layout representations that scale more naturally, such as graph-structured inputs or object-centric embeddings, to remove these spatial limitations.
Second, while OGC evaluates coordination under environmental and partner variation, it does not explicitly test agents’ ability to reason about their partners’ beliefs, intentions, or mental models. Such capabilities – often studied under theory-of-mind or agent modelling frameworks (Rabinowitz et al., 2018; Bard et al., 2020; Gandhi et al., 2021; Bara et al., 2023; Bortoletto et al., 2024b; a) – are likely to be important for achieving robust zero-shot human-AI collaboration. Future work could explore reasoning about other agents in previously unexplored environments.
Finally, while our experiments focus on DCD methods, OGC is compatible with a wide range of generalisation approaches. Future work could for instance explore few-shot adaption to partners and environments (Bauer et al., 2023) or explore representation learning approaches for better generalisation. We hope that the benchmark we provide will inspire and support further work in these areas.
@@ 8 Conclusion
We introduced the Overcooked Generalisation Challenge (OGC), the first open-source benchmark for evaluating cooperative multi-agent reinforcement learning (MARL) agents on both environment and partner generalisation. Built on Overcooked-AI and integrated with dual curriculum design (DCD) methods, OGC enables procedural training and rigorous testing in complex, multi-object environments. Compared to prior UED benchmarks, OGC presents a significantly larger and more structured design space, exposing key limitations in existing environment generation and coordination strategies. Through extensive experiments, we demonstrated that even state-of-the-art DCD algorithms struggle to train agents that generalise across layouts and partners. These findings position OGC as a diagnostic tool for probing the frontiers of generalisable cooperation. Beyond DCD research, OGC also provides infrastructure for evaluating human-AI interaction through ad-hoc teamwork and zero-shot coordination. We hope that OGC will catalyse the development of new learning methods – what we denote UED-ZSC algorithms – that jointly address the challenges of task and partner diversity in open-ended multi-agent settings.
Broader Impact Statement
This work introduces a benchmark for studying generalisation in cooperative multi-agent learning. While fundamental, it may inform future systems that support human-AI collaboration in domains such as assistive robotics or simulation-based training. Poor generalisation in such settings could lead to coordination breakdowns if deployed without safeguards.
Acknowledgments
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer. A. Penzkofer was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016.
@@ A.2 Infrastructure & Tools
We ran our experiments on a server running Ubuntu 22.04, equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs. All training runs are executed on a single GPU only. We trained our models using Jax (Bradbury et al., 2018) and Flax (Heek et al., 2023) with 1, 2, 3, 4 and 5 as random seed for training DCD methods and 1 to 8 as random seeds for the populations. Training the DCD methods usually finishes in under 24 hours, only SoftMoE and PAIRED-based methods take longer. SoftMoE-based policies often take an extra 50% wall-clock time to train. Noticeable is also that our S5 implementation is the fastest, usually needing roughly 30% less time. Both are compared to the default architectures’ training time. In the longest case, the combination of a SoftMoE-LSTM policy trained with PAIRED takes about 80 hours to complete training. Our benchmark should be runnable on any system that features a single CUDA-compatible GPU. Although in our experience our experiments will require 32GB VRAM to run.
@@ A.3 Extended Related Work
We present an overview over how our environment compares to other UED environments in Table 3.
@@ A.4 Hyperparameters
Learning Rate $\eta$ $1\cdot 10^{-4}$, $\mathbf{3\cdot 10^{-4}}$, $5\cdot 10^{-4}$, $1\cdot 10^{-3}$]
Entropy Coefficient $\mathbf{0.01}$ $0.1$]
@@
# Hidden Layers [22, $\mathbf{3}$, $4$]
Reward Shaping Annealing Steps [0, $2500000$, $5000000$, until end]
@@
$n$ walls to place Sampled between 01515
$n$ onion piles to place Sampled uniformly between 1122
$n$ plate piles to place Sampled uniformly between 1122
$n$ pots to place Sampled uniformly between 1122
$n$ goals to place Sampled uniformly between 1122
@@
$n$ walls to place Sampled between 01515
@@
In the experiment displayed in Figure 7 we show how policies behave when trained on all versus on only one layout and then decide on the SoftMoE architecture for our agents. Since these are considerably easier problems we train these models on fewer environment steps in total. We train the overfitting baseline for roughly $1/30$ ($12,800,000$ steps in the environment) of the experience of all UED methods and only use $1,000$ outer loops. For the Oracle baseline, we use $1/2$ of the experience ($200,000,000$ steps in the environment) and only $5,000$ outer loops. Both are trained until convergence, and to speed up training, we deploy 100 environment simulators. Recall that the UED methods use $30,000$ training loops and $400,000,000$ steps in the environment. Notably, in the case of population PAIRED these steps apply per student and do not include any additional steps taken in the teacher environment.
@@ A.6.2 Performance Across Levels
Table 12: Mean Solve rate averaged over the respective testing layouts. The best result is shown in bold. We report aggregate statistics over five random seeds. As a baseline we include an Oracle version for all architectures, which was trained on the five testing layouts directly.
@@
DR $4.81\pm 13.3\%$ $3.78\pm 9.8\%$ $0.00\pm 0.0\%$
PLR⟂,∥ $0.03\pm 0.0\%$ $3.68\pm 9.9\%$ $0.00\pm 0.0\%$
Pop. PAIRED $0.08\pm 0.2\%$ $\mathbf{14.72\pm 12.2\%}$ $0.00\pm 0.0\%$
ACCEL $00\pm 0.0\% 0.0$$4\pm 0.1\% 0.0$$00\pm 0.0\%$
Oracle $7.2\pm 7.7\% 99$$9.7\pm 0.6\%$ $97.2\pm 2.8\%$
To accompany the overall performance measured by reward in the main paper in Table 1 we also measure the mean solved rate in Table 12.
@@ A.6.3 Performance on Individual Levels
Table 13: Performance on all evaluation layouts. We show the mean episode reward R and the mean episode solved rate SR. The overall best result per layout is presented in bold, excluding oracle results.
@@
Cramped DR 0.02 0.0% 0.01 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.16 2.1% 0.01 0.0%
Pop. PAIRED 0.78 0.0% 8.92 9.0% 0.00 0.0%
ACCEL 1.16 2.7% 1.52 3.4% 0.28 2.1%
Oracle 197.56 100.0% 201.56 100.0% 129.64 94.0%
Coord DR 4.98 0.0% 13.2 0.0% 0.00 0.0%
PLR⟂,∥ 0.09 0.0% 0.44 6.6% 0.06 0.0%
Pop. PAIRED 0.4 0.0% 28.76 18.0% 0.1 0.0%
ACCEL 0.00 0.0% 0.28 0.0% 0.00 0.0%
Oracle 278.56 100.0% 293.00 100.0% 252.6 100.0%
Forced DR 1.78 0.0% 3.48 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.50 0.5% 0.00 0.0%
Pop. PAIRED 0.04 0.0% 2.64 6.0% 0.00 0.0%
ACCEL 0.00 0.0% 0.04 0.0 % 0.00 0.0 %
Oracle 194.12 100.0% 204.96 100.0% 146.69 96.0%
Asymm DR 7.32 0.1% 3.38 4.4% 0.32 0.0%
PLR⟂,∥ 0.82 0.0% 3.29 11.2% 0.62 0.1%
Pop. PAIRED 1.68 0.6% 33.42 40.4% 0.10 0.0%
ACCEL 0.04 0.0% 1.20 0.0 % 0.00 0.0 %
Oracle 238.68 100.0% 242.40 98.4% 206.36 99.7%
Counter DR 0.11 0.0% 0.96 0.0% 0.00 0.0%
@@
Oracle 113.28 86.0% 142.00 100.0% 95.60 96.0%
@@ A.6.4 Population Training Details
Both populations were trained over 8 random seeds. As architecture, we used a simple CNN encoder without RNN as in prior work Zhao et al. (2023); Yu et al. (2023). To give an intuition into the performance of the members of the population, we present the training curves over all 8 seeds of training an FCP population in Figure 13. MEP was trained with the same architecture, with the same amount of experience per agent and achieved practically identical results. As in prior work (Zhao et al., 2023) we set the population entropy coefficient during training to $\alpha=0.01$.
@@ A.6.5 Detailed Results with Populations
DR $19.47\pm 8.79$ $9.02\pm 4.67$ $59.62\pm 10.49$ $1.32\pm 0.60$ $21.63\pm 11.01$
PLR⟂,∥ $8.75\pm 1.99$ $7.00\pm 2.13$ $60.85\pm 5.03$ $1.78\pm 0.75$ $20.28\pm 2.41$
Pop. PAIRED $\mathbf{45.71\pm 14.65}$ $\mathbf{11.72\pm 2.49}$ $\mathbf{72.2\pm 13.86}$ $\mathbf{2.46\pm 1.05}$ $\mathbf{23.99\pm 4.94}$
ACCEL $11.30\pm 6.37$ $10.03\pm 2.43$ $59.83\pm 4.43$ $2.46\pm 1.05$ $18.29\pm 1.69$
@@
DR $21.18\pm 10.27$ $18.93\pm 4.26$ $56.63\pm 10.28$ $1.86\pm 0.74$ $32.19\pm 6.23$
PLR⟂,∥ $8.89\pm 2.03$ $20.68\pm 1.61$ $56.66\pm 3.38$ $2.38\pm 1.13$ $29.71\pm 2.24$
Pop. PAIRED $\mathbf{39.02\pm 22.74}$ $\mathbf{20.85\pm 5.72}$ $\mathbf{64.51\pm 8.74}$ $\mathbf{2.95\pm 2.04}$ $\mathbf{33.29\pm 6.11}$
ACCEL $12.00\pm 6.40$ $19.06\pm 1.73$ $51.94\pm 5.28$ $2.42\pm 2.44$ $28.91\pm 1.51$
@@
DR $25.99\pm 13.37\%$ $6.49\pm 4.65\%$ $67.85\pm 8.57\%$ $0.03\pm 0.05\%$ $30.26\pm 21.34\%$
PLR⟂,∥ $10.95\pm 3.10\%$ $3.82\pm 2.18\%$ $68.11\pm 1.83\%$ $0.07\pm 0.05\%$ $25.99\pm 7.39\%$
Pop. PAIRED $\mathbf{60.18\pm 18.29\%}$ $\mathbf{9.08\pm 4.51\%}$ $\mathbf{77.71\pm 83.7\%}$ $0.21\pm 0.30\%$ $\mathbf{35.16\pm 11.97\%}$
ACCEL $13.84\pm 10.17\%$ $7.11\pm 3.77\%$ $67.26\pm 2.48\%$ $\mathbf{0.31\pm 0.68\%}$ $21.33\pm 3.81\%$
@@
DR $29.51\pm 17.44\%$ $28.18\pm 7.52\%$ $65.81\pm 8.62\%$ $0.10\pm 0.09\%$ $51.61\pm 6.33\%$
PLR⟂,∥ $10.26\pm 3.43\%$ $\mathbf{31.81\pm 4.53\%}$ $65.71\pm 3.43\%$ $0.26\pm 0.32\%$ $\mathbf{49.97\pm 3.64\%}$
Pop. PAIRED $\mathbf{50.39\pm 29.73\%}$ $30.66\pm 12.23\%$ $\mathbf{74.21\pm 7.05\%}$ $0.29\pm 0.49\%$ $45.16\pm 17.49\%$
ACCEL $14.16\pm 10.53\%$ $27.36\pm 4.51\%$ $63.61\pm 4.46\%$ $\mathbf{0.83\pm 1.81\%}$ $49.71\pm 2.32\%$
@@
A.7 Training Curves and Evaluation